If the point (3,10) is on the graph of y=a^x, what is this point on the graph of y=log(base)a X
I am confused and need help. I dont even know where to start. Can you help?

Answer by lwsshak3(6498)

If the point (3,10) is on the graph of y=a^x, what is this point on the graph of y=log(base)a X
**
I'm not absolutely sure, but I think this is the answer:
y=loga(x) is the logarithm form
and
y=a^(x) is the exponential form
The two forms are inverses of each other.
If I remember correctly, the domain and range switch between them, that is the domain and range of one form is the range and domain of the other form.
So, (3,10) on y=a^(x) appears as (10,3) on y=loga(x)
If you graph the two functions you can see that (0,1) is on the exponential curve and (1,0) is on the log curve